OpenStudy (anonymous):

Help!!! Trigonometric Substitution! Integral (x/sqrt(x^2 - 4x + 8)

OpenStudy (anonymous):

\[\int\limits \frac{ x }{ \sqrt{x^{2}-4x+8} }dx\]

OpenStudy (anonymous):

I know that, \[x=2\tan \theta + 2\]

OpenStudy (anonymous):

\[dx=2\sec^2 \theta d \theta \]

zepdrix (zepdrix):

I'm curious how you're approaching this. I would start by completing the square in the denominator. Makes it easier to see the substitution. \[\Large \int\limits \frac{x}{\sqrt{(x-2)^2+4}}dx\]

OpenStudy (anonymous):

Yeah! I did that, too!

zepdrix (zepdrix):

ok cool :)

zepdrix (zepdrix):

Hmm yah your sub looks good so far. Just having trouble plugging everything in or what? :o

OpenStudy (anonymous):

Yeah, that's my problem!

zepdrix (zepdrix):

Let's get this last little piece of information:\[\Large x=2\tan \theta+2 \qquad\to\qquad (x-2)=2\tan \theta\] I only placed the brackets there so it's easier to see where we're going to place it in the integral.

zepdrix (zepdrix):

\[\large \int\limits\limits \frac{x}{\sqrt{(x-2)^2+4}}dx \qquad\to\qquad \int\limits\limits \frac{2\tan \theta+2}{\sqrt{(2\tan \theta)^2+4}}(2\sec^2\theta\;d \theta)\]

zepdrix (zepdrix):

Understand how those are plugging in? :o We had 3 pieces of information we had to use to make that substitution.

OpenStudy (anonymous):

Yeah, so far, so good! :-)

zepdrix (zepdrix):

\[\Large \int\limits \frac{2\tan \theta+2}{\sqrt{4\tan^2\theta+4}}(2\sec^2\theta \;d \theta)\]Hmm looks like we'll have some nice cancellations from here.

OpenStudy (anonymous):

Okay, you cancel out \[4\tan^2 \theta \]

OpenStudy (anonymous):

with \[2\sec^2 \theta\]

OpenStudy (anonymous):

Right?

zepdrix (zepdrix):

Sorry for the delay :( OpenStudy crashing again.. grr \[\large \int\limits\limits \frac{2\tan \theta+2}{\sqrt{4(\tan^2\theta+1)}}(2\sec^2\theta \;d \theta) \quad=\quad \int\limits\limits \frac{2\tan \theta+2}{\sqrt{4(\sec^2\theta)}}(2\sec^2\theta \;d \theta)\]\[\large =\int\limits \frac{2\tan \theta+2}{2\sec \theta}(2\sec^2\theta\;d \theta)\]

zepdrix (zepdrix):

So it looks like it will only cancel out with one of the secants up top right? Because we had to take the square root in the bottom, losing a power.

OpenStudy (anonymous):

Yes! I think so, too!

zepdrix (zepdrix):

Mmm ok that simplified down a bit for us!\[\Large \int\limits (2 \tan \theta+2)(\sec \theta\; d \theta)\]

OpenStudy (anonymous):

Don't we have to do the right triangle?

zepdrix (zepdrix):

At some point, yes we'll need to do that. Whether now or later, it doesn't matter.

zepdrix (zepdrix):

I think this is the relationship we'll want to draw, yes?\[\Large \tan \theta= \frac{x-2}{2}\]

zepdrix (zepdrix):

Understand how to label this triangle? :)

OpenStudy (anonymous):

I do! But for this problem in specific, I'm not sure how I will label it

zepdrix (zepdrix):

\[\Large x=2\tan \theta+2 \qquad\to\qquad \tan \theta=\frac{x-2}{2}=\frac{opposite}{adjacent}\]

OpenStudy (anonymous):

Go it! So the will the hypotenuse be \[\sqrt{x^2-4x+8}\]

OpenStudy (anonymous):

?

zepdrix (zepdrix):

ya looks good.

OpenStudy (anonymous):

Okay, now, how do you get to the final answer?

zepdrix (zepdrix):

We don't want to use the triangle just yet though! :O Our integral is written in terms of trig functions (which will make the integration step much easier for us). After we integrate, we can undo our substitutions by using the triangle.

zepdrix (zepdrix):

\[\Large \int\limits\limits (2 \tan \theta+2)(\sec \theta\; d \theta) \quad=\quad 2\int\limits \sec \theta \tan \theta+ \sec \theta \;d \theta\]

zepdrix (zepdrix):

Remember how to integrate those two terms or no? :D

OpenStudy (anonymous):

I would be great if we can go over them again! I just want to make sure that I fully understand the concept!

OpenStudy (anonymous):

It would*

zepdrix (zepdrix):

There's not much to the first term, just something to remember.\[\Large (\sec \theta)' \quad=\quad \sec \theta \tan \theta\] Does that look familiar? :o

OpenStudy (anonymous):

Yes, it does!

zepdrix (zepdrix):

Again for this other term, it's one that you'll simply want to memorize:\[\Large \int\limits \sec \theta\; d \theta \qquad=\qquad \ln\left|\sec \theta+\tan \theta\right|\]We can go through the steps if you want, but there's not a lot of intuition in solving it.

OpenStudy (anonymous):

Okay, got it!

OpenStudy (anonymous):

When do you we use the right triangle in order to arrive to the final answer?

OpenStudy (anonymous):

The answer is supposed to be \[2\ln(\sqrt{x^2-4x+8}+x-2)+ \sqrt{x^2-4x+8}\] \]

zepdrix (zepdrix):

So we integrated getting:\[\large 2\int\limits \sec \theta \tan \theta + \sec \theta\; d \theta \qquad=\qquad 2\left[\sec \theta+\ln\left(\sec \theta+\tan \theta\right)\right]+C\]

zepdrix (zepdrix):

Now we can finally use our triangle.

zepdrix (zepdrix):

We already have tangent, right? \(\Large \dfrac{x-2}{2}\) We'll also need to figure out secant from the triangle we set up.

zepdrix (zepdrix):

Still confused on that integration step? :3

OpenStudy (anonymous):

No, I got it!

OpenStudy (anonymous):

Proceed

zepdrix (zepdrix):

based on our triangle we drew, it looks like:\[\large \sec \theta\quad=\quad \frac{\sqrt{x^2-4x+8}}{2}\]Hmm the square root bar is tweaking out on me -_- should be above the entire numerator there.

OpenStudy (anonymous):

Okay!

zepdrix (zepdrix):

So we just need to use those relationships to undo our substitutions now:\[\Large 2\color{#CC0033}{\sec \theta}+2\ln\left(\color{#CC0033}{\sec \theta}+\color{#3366CF}{\tan \theta}\right)+C\]

zepdrix (zepdrix):

\[\Large 2\color{#CC0033}{\frac{\sqrt{x^2-4x+8}}{2}}+2\ln\left(\color{#CC0033}{\frac{\sqrt{x^2-4x+8}}{2}}+\color{#3366CF}{\tan \theta}\right)+C\]

zepdrix (zepdrix):

Understand how to replace the tangent also?

OpenStudy (anonymous):

I do! So, the 2's cancel out, and the final answer is \[2\ln (\sqrt{x^2-4x+8}+x-2)+\sqrt{x^2-4x+8}\]

OpenStudy (anonymous):

Rigth?

zepdrix (zepdrix):

yah the 2's cancel out in that term. Where did the 2 inside the log go? :o hmm

zepdrix (zepdrix):

We should have:\[\large 2\ln\left(\frac{\sqrt{x^2-4x+8}+x-2}{2}\right)+\sqrt{x^2-4x+8}+C\]

zepdrix (zepdrix):

To match the answer in the book or whatever, we would need to apply another step. We need to use a rule of logarithms.\[\large \log\left(\frac{a}{b}\right) \quad=\quad \log(a)-\log(b)\]

OpenStudy (anonymous):

Oh, okay! But the answer we got is correct, right?

zepdrix (zepdrix):

\[2\ln\left(\sqrt{x^2-4x+8}+x-2\right)\color{orangered}{-2\ln(2)}+\sqrt{x^2-4x+8}+\color{orangered}{C}\] From here 2ln(2) is just a constant so we can absorb it into the C.

zepdrix (zepdrix):

Yah our answer was correct :) No need to go any further unless you want to.

zepdrix (zepdrix):

Makes it a little tricky when the books oversimplify these things lol

OpenStudy (anonymous):

Oh wow! Really, THANK YOU SO MUCH for all the time you spend! I sincerely appreciate it! You were really clear and had patient! I now understand! THANK YOU! :-)

zepdrix (zepdrix):

yay team \c:/ np!